# Enhancing Variable Selection in Large-scale Logistic Regression: Leveraging Manual Labeling with Beneficial Noise

### v1

@@ A-C1 Basic Lemma and Four Matrices
$\bm{Q}=\begin{pmatrix}\bm{E}_{1}&\bm{0}&\bm{0}\\ \bm{0}&\mu\bm{I}_{n}+\bm{E}_{2}&\bm{0}\\ \bm{0}&\bm{I}_{n}&\frac{1}{\mu}\bm{I}_{n}\end{pmatrix},$
$\bm{M}=\begin{pmatrix}\bm{I}_{p}&\bm{0}&\bm{0}\\ \bm{0}&\bm{I}_{n}&\bm{0}\\ \bm{0}&\mu\bm{I}_{n}&\bm{I}_{n}\end{pmatrix},$ (51)
$\bm{H}=\begin{pmatrix}\bm{E}_{1}&\bm{0}&\bm{0}\\ \bm{0}&\mu\bm{I}_{n}+\bm{E}_{2}&\bm{0}\\ \bm{0}&\bm{0}&\frac{1}{\mu}\bm{I}_{n}\end{pmatrix},$
$\bm{G}=\begin{pmatrix}\bm{E}_{1}&\bm{0}&\bm{0}\\ \bm{0}&\bm{E}_{2}&\bm{0}\\ \bm{0}&\bm{0}&\frac{1}{\mu}\bm{I}_{n}\end{pmatrix}.$ (52)
@@
$\bm{HM}=\bm{Q}\ \text{and}\ \bm{G}=\bm{Q}^{\top}+\bm{Q}-\bm{M}^{\top}\bm{H}\bm{M}.$ (53)
A-C2 Variational inequality characterization
@@
$\min_{\bm{\beta},\bm{r}}$ $L(\bm{r})+\lambda\|\bm{w}\odot\bm{\beta}\|_{1},$ (54)
@@
where $L(\bm{r})=\sum_{i=1}^{n}\left[-S_{i}r_{i}+m\log\left(1+\exp({r_{i}})\right)\right]$. And the Lagrange multiplier form of (54) is
$L(\bm{\beta},\bm{r};\bm{u})=L(\bm{r})+\lambda\|\bm{w}\odot\bm{\beta}\|_{1}-\bm{u}^{\top}(\bm{X}\bm{\beta}-\bm{r}).$ (55)
According to the variational inequality characterization in section 2 of [41], we know that the solution of the constrained optimization function above is the saddle point of the following Lagrangian function in (55). As described in [41], finding a saddle point of $L(\bm{\beta},\bm{r};\bm{u})$ is equivalent to finding $\bm{\beta}^{*},\bm{r}^{*},\bm{u}^{*}$ such that the following inequalities are satisfied:
$\bm{v}^{*}\in\Omega,\ L(\bm{r})+\lambda\|\bm{w}\odot\bm{\beta}\|_{1}-L(\bm{r}^{*})-\lambda\|\bm{w}\odot\bm{\beta}^{*}\|_{1}$
$+(\bm{v}-\bm{v}^{*})^{\top}F(\bm{v}^{*})\geq 0,\ \forall\bm{v}\in\Omega,$ (56)
@@
$F(\bm{v})=\begin{pmatrix}-\bm{X}^{\top}\bm{u}\\ \bm{u}\\ \bm{X}\bm{\beta}-\bm{r}\end{pmatrix}.$ (57)
Note that the operator $F$ defined in (57) is monotone, because
$(\bm{v}_{1}-\bm{v}_{2})^{\top}\left[F(\bm{v}_{1})-F(\bm{v}_{2})\right]\equiv 0,\quad\forall\ \bm{v}_{1},\bm{v}_{2}\in\Omega.$ (58)
Throughout, we denote by $\Omega^{*}$ the solution set of (A-C2), which is also the set of saddle points of the Lagrangian function (55) of the model (54).
A-C3 Proof
@@
{${l}\check{\bm{\beta}}^{k}&\leftarrow\mathop{\arg\min}\limits_{\bm{\beta}}\left\{\lambda\|\bm{w}\odot\bm{\beta}\|_{1}+\frac{\mu}{2}\|\bm{X}\bm{\beta}-\bm{r}^{k}\right.\\ &\left.\qquad-\bm{u}^{k}/\mu\|_{2}^{2}+\frac{1}{2}\|\bm{\beta}-\bm{\beta}^{k}\|^{2}_{{\bm{E}}_{1}}\right\};\\ \check{\bm{r}}^{k}&\leftarrow\mathop{\arg\min}\limits_{\bm{r}}\left\{L(\bm{r})+\frac{\mu}{2}\|\bm{X}\bm{\check{}}{\bm{\beta}}^{k}-\bm{r}-\bm{u}^{k}/\mu\|_{2}^{2}\right.\\ &\left.\qquad+\frac{1}{2}\|\bm{r}-\bm{r}^{k}\|^{2}_{\bm{E}_{2}}\right\};\\ \check{\bm{u}}^{k}&\leftarrow\bm{u}^{k}-\mu(\bm{X}\check{\bm{\beta}}^{k}-{\bm{r}}^{k}),\end{aligned}\right.$ (59)
The above iterative steps can be written in a form similar to (A-C2), and the specific cases can be summarized in the following lemma.
@@
Given $\bm{v}^{k}$, for any $\bm{v}\in\Omega$, the predicted variable $\check{\bm{v}}^{k}$ generated by (59) satisfies
$\ L(\bm{r})+\lambda\|\bm{w}\odot\bm{\beta}\|_{1}-L(\check{\bm{r}}^{k})-\lambda\|\bm{w}\odot\check{\bm{\beta}}^{k}\|_{1}$
$+$ $(\bm{v}-\check{\bm{v}}^{k})^{\top}F(\check{\bm{v}}^{k})\geq(\bm{v}-\check{\bm{v}}^{k})^{\top}\bm{Q}({\bm{v}}^{k}-\check{\bm{v}}^{k}),$ (60)
where $\bm{Q}$ is defined in (A-C1).
Proof:
According to Lemma 2 of the first-order optimality principle, the optimality condition for the subproblem of $\check{\bm{\beta}}^{k}$ in (59) is
$\lambda\|\bm{w}\odot\bm{\beta}\|_{1}-\lambda\|\bm{w}\odot\check{\bm{\beta}}^{k}\|_{1}$
+\ + $(\bm{\beta}-\check{\bm{\beta}}^{k})^{\top}\left[\mu\bm{X}^{\top}(\bm{X}\check{\bm{\beta}}^{k}-\bm{r}^{k}-\bm{u}^{k}/\mu)\right.$
+\ + $\left.\bm{E}_{1}(\check{\bm{\beta}}^{k}-{\bm{\beta}}^{k})\right]\geq 0.$
Using $\check{\bm{u}}^{k}=\bm{u}^{k}-\mu(\bm{X}\check{\bm{\beta}}^{k}-{\bm{r}}^{k})$ in (59), the above equation is
$\lambda\|\bm{w}\odot\bm{\beta}\|_{1}-\lambda\|\bm{w}\odot\check{\bm{\beta}}^{k}\|_{1}$ (61)
+\ + $(\bm{\beta}-\check{\bm{\beta}}^{k})^{\top}(-\bm{X}^{\top}\check{\bm{u}}^{k}+\bm{E}_{1}(\check{\bm{\beta}}^{k}-{\bm{\beta}}^{k}))\geq 0.$
@@
$L(\bm{r})-L(\check{\bm{r}}^{k})$ (62)
$+$ $(\bm{r}-\check{\bm{r}}^{k})^{\top}\left[\check{\bm{u}}^{k}+\mu(\check{\bm{r}}^{k}-\bm{r}^{k})+\bm{E}_{2}(\check{\bm{r}}^{k}-\bm{r}^{k})\right]\geq 0.$
The formula for generating $\check{\bm{u}}^{k}$ in (59) can also be written as
@@
$(\bm{u}-\check{\bm{u}}^{k})^{\top}\left[(\bm{X}\check{\bm{\beta}}^{k}-\check{\bm{r}}^{k})+(\check{\bm{r}}^{k}-{\bm{r}}^{k})+\frac{1}{\mu}(\check{\bm{u}}^{k}-{\bm{u}}^{k})\right]\geq 0.$ (63)
Combining (61), (62), and (63) together can lead to the conclusion of Lemma (3).∎
Note that the iteration steps of (59) are the same as the iteration variables $\bm{\beta}$ and $\bm{r}$ generated by our proposed Algorithm 1, but the iteration of $\bm{u}$ is different. Therefore, we need the following correction step,
$\bm{v}^{k+1}=\bm{v}^{k}-\bm{M}(\bm{v}^{k}-\check{\bm{v}}^{k}),$ (64)
where $\bm{M}$ is defined in (A-C1). The above discussion shows that the proposed Algorithm 1 can be divided into two steps: prediction in (59) and correction in (64). This prediction and correction step has been used as an effective tool to demonstrate the convergence of the ADMM algorithm, as detailed in [40, 42] and its references.
@@
It can be easily inferred from (A-C1) and (53) that Algorithm 1 satisfies the above convergence conditions. Then, from the results of Theorems 5.2 and 5.6 of [40], we have
$\|\bm{v}^{k+1}-\bm{v}^{*}\|_{\bm{H}}^{2}\leq\|\bm{v}^{k}-\bm{v}^{*}\|_{\bm{H}}^{2}-\|\bm{v}^{k}-\check{\bm{v}}^{k}\|_{\bm{G}}^{2},$ (65)
$\|\bm{M}(\bm{v}^{k+1}-\check{\bm{v}}^{k+1})\|_{\bm{H}}^{2}\leq\|\bm{M}(\bm{v}^{k}-\check{\bm{v}}^{k})\|_{\bm{H}}^{2}.$ (66)
The inequality in (65) implies that the sequence $\{\bm{v}^{k}\}$ exhibits Fejér monotonicity. When $\bm{G}\succ 0$, this sequence converges to a point $\bm{v}^{*}\in\Omega^{*}$ in the $\bm{H}$-norm. This key inequality is fundamental for establishing the global convergence of the sequence. The proof of the global convergence can be found in Theorem 3.2 of [42]. So far, the proof of the first property of Theorem 3 has been completed.
Since $\bm{M}(\bm{v}^{k+1}-\check{\bm{v}}^{k+1})=\bm{v}^{k+1}-\bm{v}^{k+2}$ and $\bm{M}(\bm{v}^{k}-\check{\bm{v}}^{k})=\bm{v}^{k}-\bm{v}^{k+1}$, the inequality in (66) implies that
$\|\bm{v}^{k+1}-\bm{v}^{k+2}\|_{\bm{H}}^{2}\leq\|\bm{v}^{k}-\bm{v}^{k+1}\|_{\bm{H}}^{2}.$ (67)
From the inequality in (65), it follows from $\bm{G}\succ 0$, there is a constant $c_{0}>0$ such that
$\|\bm{v}^{k+1}-\bm{v}^{*}\|_{\bm{H}}^{2}$ (68)
$\leq$ $\|\bm{v}^{k}-\bm{v}^{*}\|_{\bm{H}}^{2}-c_{0}\|\bm{M}(\bm{v}^{k}-\check{\bm{v}}^{k})\|_{\bm{H}}^{2}$
$=$ $\|\bm{v}^{k}-\bm{v}^{*}\|_{\bm{H}}^{2}-c_{0}\|\bm{v}^{k}-\bm{v}^{k+1}\|_{\bm{H}}^{2}.$
By summing up inequality in (68) over $k=0,1,\cdots,K$, we obtain
@@
Combining (67) leads to the conclusion that
$c_{0}(K+1)\|\bm{v}^{K}-\bm{v}^{K+1}\|_{\bm{H}}^{2}\leq\|\bm{v}^{0}-\bm{v}^{*}\|_{\bm{H}}^{2}.$ (69)
@@ A-D Proof of Throrem 4
$\left\{\hat{\bm{\beta}}^{k},\hat{\bm{z}}^{k},\hat{\bm{u}}^{k}\right\}=\left\{\bar{\bm{\beta}}^{k},\bar{\bm{z}}^{k},\bar{\bm{u}}^{t}\right\}.$ (70)
@@
$\hat{\bm{\beta}}^{k+1}\leftarrow$ $\arg\min_{\bm{\beta}}\left\{\lambda\|\bm{w}\odot\bm{\beta}\|_{1}\right.$ (71)
$+$ $\left.\frac{\eta}{2}\|\bm{\beta}-\hat{\bm{\beta}}^{k}+\frac{\mu\bm{X}^{\top}(\bm{X}\hat{\bm{\beta}}^{k}-\hat{\bm{r}}^{k}-\hat{\bm{u}}^{k}/\mu)}{\eta}\|_{2}^{2}\right\}.$
@@
$\bar{\bm{\beta}}^{k+1}\leftarrow\underset{\bm{\beta}}{\mathrm{argmin}}$$\left\{\lambda\left\lVert\bm{w}\odot\bm{\beta}\right\rVert_{1}\right.$ (72)
$+$ $\left.\frac{\eta}{2}\left\lVert\bm{\beta}-\bar{\bm{\beta}}^{k}+\sum_{g=1}^{G}\bm{\xi}_{g}^{k}\right\rVert_{2}^{2}\right\},$
@@
$\sum_{g=1}^{G}\bm{\xi}_{g}^{k}=\frac{\mu\bm{X}^{\top}(\bm{X}\bar{\bm{\beta}}^{k}-\bar{\bm{r}}^{k}-\bar{\bm{u}}^{k}/\mu)}{\eta}.$ (73)
According to the assumption in (70), there is $\hat{\bm{\beta}}^{k}=\bar{\bm{\beta}}^{k}$, $\hat{\bm{r}}^{k}=\bar{\bm{r}}^{k}$ and $\hat{\bm{u}}^{k}=\bar{\bm{u}}^{k}$. It follows from (71) and (72) that
$\hat{\bm{\beta}}^{k+1}=\bar{\bm{\beta}}^{k+1}.$ (74)
@@
$\hat{\bm{r}}^{k+1}\leftarrow$ $\left[\Phi(\hat{\bm{r}}^{k})+\mu\bm{I}_{p}\right]^{-1}\left[\mu(\bm{X}\hat{\bm{\beta}}^{k+1}-\hat{\bm{u}}^{k}/\mu)\right.$ (75)
$+$ $\left.\Phi(\hat{\bm{r}}^{k})\hat{\bm{r}}^{k}-\phi^{{}^{\prime}}(\hat{\bm{r}}^{k})+\bm{S}\right].$
$\hat{\bm{r}}^{k+1}$ can be expressed as $[(\hat{\bm{r}}_{1}^{k+1})^{\top},(\hat{\bm{r}}_{2}^{k+1})^{\top},\ldots,(\hat{\bm{r}}_{G}^{k+1})^{\top}]$, and $\bm{S}$, $\hat{\bm{u}}^{k}$ and $\hat{\bm{r}}^{k}$ are divided in a similar manner. Drawing on the definition of $\Phi(\hat{\bm{r}}^{k})$, it becomes clear that the diagonal elements of its block-diagonal matrix are $\Phi(\hat{\bm{r}}_{g}^{k})$, with $g$ ranging from $1$ to $G$. Then, we have
$\hat{\bm{r}}_{g}^{k+1}\leftarrow$ $\left[\Phi(\hat{\bm{r}}_{g}^{k})+\mu\bm{I}_{n_{g}}\right]^{-1}\left[\mu(\bm{X}\hat{\bm{\beta}}_{g}^{k+1}-\hat{\bm{u}}_{g}^{k}/\mu)\right.$ (76)
$+$ $\left.\Phi(\hat{\bm{r}}_{g}^{k})\hat{\bm{r}}_{g}^{k}-\phi^{{}^{\prime}}(\hat{\bm{r}}_{g}^{k})+\bm{S}_{g}\right],g=1,\dots,G.$
@@
$\bar{\bm{r}}_{g}^{k+1}\leftarrow$ $\left[\Phi(\bar{\bm{r}}_{g}^{k})+\mu\bm{I}_{n_{g}}\right]^{-1}\left[\mu(\bm{X}_{g}\bar{\bm{\beta}}^{k+1}-\bar{\bm{u}}_{g}^{k}/\mu)\right.$
$+$ $\left.\Phi(\bar{\bm{r}}_{g}^{k})\bar{\bm{r}}_{g}^{k}-\phi^{{}^{\prime}}(\bar{\bm{r}}_{g}^{k})+\bm{S}_{g}\right],g=1,\dots,G.$ (77)
Again, according to the assumption in (70) and (74), there is $\hat{\bm{r}}_{g}^{k}=\bar{\bm{r}}_{g}^{k}$, $\hat{\bm{u}}_{g}^{k}=\bar{\bm{u}}_{g}^{k}$ and $\hat{\bm{\beta}}^{k+1}=\bar{\bm{\beta}}^{k+1}$. It follows from (76) and (A-D) that
$\hat{\bm{r}}_{g}^{k+1}=\bar{\bm{r}}_{g}^{k+1},g=1,\dots,G.$ (78)
@@
It follows from (70), (74) and (78) that
$\hat{\bm{u}}_{g}^{t+1}=\bar{\bm{u}}_{g}^{t+1},g=1,\dots,G.$ (81)
@@
$\left\{\hat{\bm{\beta}}^{k},\hat{\bm{r}}^{k},\hat{\bm{u}}^{k}\right\}=\left\{\bar{\bm{\beta}}^{k},\bar{\bm{z}}^{k},\bar{\bm{u}}^{k}\right\},\quad\text{for all }k.$ (83)

### v2

@@ C.1 Basic Lemma and Four Matrices
$\bm{Q}=\begin{pmatrix}\bm{E}_{1}&\bm{0}&\bm{0}\\ \bm{0}&\mu\bm{I}_{n}+\bm{E}_{2}&\bm{0}\\ \bm{0}&\bm{I}_{n}&\frac{1}{\mu}\bm{I}_{n}\end{pmatrix},\ \ \bm{M}=\begin{pmatrix}\bm{I}_{p}&\bm{0}&\bm{0}\\ \bm{0}&\bm{I}_{n}&\bm{0}\\ \bm{0}&\mu\bm{I}_{n}&\bm{I}_{n}\end{pmatrix}$, (S1.11)
$\bm{H}=\begin{pmatrix}\bm{E}_{1}&\bm{0}&\bm{0}\\ \bm{0}&\mu\bm{I}_{n}+\bm{E}_{2}&\bm{0}\\ \bm{0}&\bm{0}&\frac{1}{\mu}\bm{I}_{n}\end{pmatrix},\ \ \bm{G}=\begin{pmatrix}\bm{E}_{1}&\bm{0}&\bm{0}\\ \bm{0}&\bm{E}_{2}&\bm{0}\\ \bm{0}&\bm{0}&\frac{1}{\mu}\bm{I}_{n}\end{pmatrix}$. (S1.12)
@@
$\bm{HM}=\bm{Q}\ \text{and}\ \bm{G}=\bm{Q}^{\top}+\bm{Q}-\bm{M}^{\top}\bm{H}\bm{M}.$ (S1.13)
C.2 Variational inequality characterization
@@
$\min_{\bm{\beta},\bm{r}}$ $L(\bm{r})+\lambda\|\bm{w}\odot\bm{\beta}\|_{1},$ (S1.14)
@@
where $L(\bm{r})=\sum_{i=1}^{n}\left[-S_{i}r_{i}+m\log\left(1+\exp({r_{i}})\right)\right]$. And the Lagrange multiplier form of (S1.14) is
$L(\bm{\beta},\bm{r};\bm{u})=L(\bm{r})+\lambda\|\bm{w}\odot\bm{\beta}\|_{1}-\bm{u}^{\top}(\bm{X}\bm{\beta}-\bm{r}).$ (S1.15)
According to the variational inequality characterization in section 2 of [40], we know that the solution of the constrained optimization function above is the saddle point of the following Lagrangian function in (S1.15). As described in [40], finding a saddle point of $L(\bm{\beta},\bm{r};\bm{u})$ is equivalent to finding $\bm{\beta}^{*},\bm{r}^{*},\bm{u}^{*}$ such that the following inequalities are satisfied:
$\bm{v}^{*}\in\Omega,\ L(\bm{r})+\lambda\|\bm{w}\odot\bm{\beta}\|_{1}-L(\bm{r}^{*})-\lambda\|\bm{w}\odot\bm{\beta}^{*}\|_{1}+(\bm{v}-\bm{v}^{*})^{\top}F(\bm{v}^{*})\geq 0,\ \forall\bm{v}\in\Omega,$ (S1.16)
@@
$F(\bm{v})=\begin{pmatrix}-\bm{X}^{\top}\bm{u}\\ \bm{u}\\ \bm{X}\bm{\beta}-\bm{r}\end{pmatrix}$. (S1.17)
Note that the operator $F$ defined in (S1.17) is monotone, because
$(\bm{v}_{1}-\bm{v}_{2})^{\top}\left[F(\bm{v}_{1})-F(\bm{v}_{2})\right]\equiv 0,\quad\forall\ \bm{v}_{1},\bm{v}_{2}\in\Omega.$ (S1.18)
Throughout, we denote by $\Omega^{*}$ the solution set of (S1.16), which is also the set of saddle points of the Lagrangian function (S1.15) of the model (S1.14).
C.3 Proof
@@
$\left\{\begin{aligned} \check{\bm{\beta}}^{k}&\leftarrow\mathop{\arg\min}\limits_{\bm{\beta}}\left\{\lambda\|\bm{w}\odot\bm{\beta}\|_{1}+\frac{\mu}{2}\|\bm{X}\bm{\beta}-\bm{r}^{k}-\bm{u}^{k}/\mu\|_{2}^{2}+\frac{1}{2}\|\bm{\beta}-\bm{\beta}^{k}\|^{2}_{{\bm{E}}_{1}}\right\};\\ \check{\bm{r}}^{k}&\leftarrow\mathop{\arg\min}\limits_{\bm{r}}\left\{L(\bm{r})+\frac{\mu}{2}\|\bm{X}\bm{\check{}}{\bm{\beta}}^{k}-\bm{r}-\bm{u}^{k}/\mu\|_{2}^{2}+\frac{1}{2}\|\bm{r}-\bm{r}^{k}\|^{2}_{\bm{E}_{2}}\right\};\\ \check{\bm{u}}^{k}&\leftarrow\bm{u}^{k}-\mu(\bm{X}\check{\bm{\beta}}^{k}-{\bm{r}}^{k}),\end{aligned}\right$. (S1.19)
The above iterative steps can be written in a form similar to (S1.16), and the specific cases can be summarized in the following lemma.
@@
Given $\bm{v}^{k}$, for any $\bm{v}\in\Omega$, the predicted variable $\check{\bm{v}}^{k}$ generated by (S1.19) satisfies
$\ L(\bm{r})+\lambda\|\bm{w}\odot\bm{\beta}\|_{1}-L(\check{\bm{r}}^{k})-\lambda\|\bm{w}\odot\check{\bm{\beta}}^{k}\|_{1}+(\bm{v}-\check{\bm{v}}^{k})^{\top}F(\check{\bm{v}}^{k})\geq(\bm{v}-\check{\bm{v}}^{k})^{\top}\bm{Q}({\bm{v}}^{k}-\check{\bm{v}}^{k}),$
where $\bm{Q}$ is defined in (S1.11).
Proof. According to Lemma 2 of the first-order optimality principle, the optimality condition for the subproblem of $\check{\bm{\beta}}^{k}$ in (S1.19) is
$\lambda\|\bm{w}\odot\bm{\beta}\|_{1}-\lambda\|\bm{w}\odot\check{\bm{\beta}}^{k}\|_{1}+(\bm{\beta}-\check{\bm{\beta}}^{k})^{\top}\left[\mu\bm{X}^{\top}(\bm{X}\check{\bm{\beta}}^{k}-\bm{r}^{k}-\bm{u}^{k}/\mu)+\bm{E}_{1}(\check{\bm{\beta}}^{k}-{\bm{\beta}}^{k})\right]\geq 0.$
Using $\check{\bm{u}}^{k}=\bm{u}^{k}-\mu(\bm{X}\check{\bm{\beta}}^{k}-{\bm{r}}^{k})$ in (S1.19), the above equation is
$\lambda\|\bm{w}\odot\bm{\beta}\|_{1}-\lambda\|\bm{w}\odot\check{\bm{\beta}}^{k}\|_{1}+(\bm{\beta}-\check{\bm{\beta}}^{k})^{\top}(-\bm{X}^{\top}\check{\bm{u}}^{k}+\bm{E}_{1}(\check{\bm{\beta}}^{k}-{\bm{\beta}}^{k}))\geq 0.$ (S1.20)
@@
$L(\bm{r})-L(\check{\bm{r}}^{k})+(\bm{r}-\check{\bm{r}}^{k})^{\top}\left[\check{\bm{u}}^{k}+\mu(\check{\bm{r}}^{k}-\bm{r}^{k})+\bm{E}_{2}(\check{\bm{r}}^{k}-\bm{r}^{k})\right]\geq 0.$ (S1.21)
The formula for generating $\check{\bm{u}}^{k}$ in (S1.19) can also be written as
@@
$(\bm{u}-\check{\bm{u}}^{k})^{\top}\left[(\bm{X}\check{\bm{\beta}}^{k}-\check{\bm{r}}^{k})+(\check{\bm{r}}^{k}-{\bm{r}}^{k})+\frac{1}{\mu}(\check{\bm{u}}^{k}-{\bm{u}}^{k})\right]\geq 0.$ (S1.22)
Combining (S1.20), (S1.21), and (S1.22) together can lead to the conclusion of Lemma 3.
Note that the iteration steps of (S1.19) are the same as the iteration variables $\bm{\beta}$ and $\bm{r}$ generated by our proposed Algorithm 1, but the iteration of $\bm{u}$ is different. Therefore, we need the following correction step,
$\bm{v}^{k+1}=\bm{v}^{k}-\bm{M}(\bm{v}^{k}-\check{\bm{v}}^{k}),$ (S1.23)
where $\bm{M}$ is defined in (S1.11). The above discussion shows that the proposed Algorithm 1 can be divided into two steps: prediction in (S1.19) and correction in (S1.23). This prediction and correction step has been used as an effective tool to demonstrate the convergence of the ADMM algorithm, as detailed in [39], [41] and its references.
@@
It can be easily inferred from (S1.12) and (S1.13) that Algorithm 1 satisfies the above convergence conditions. Then, from the results of Theorems 5.2 and 5.6 of [41], we have
$\|\bm{v}^{k+1}-\bm{v}^{*}\|_{\bm{H}}^{2}\leq\|\bm{v}^{k}-\bm{v}^{*}\|_{\bm{H}}^{2}-\|\bm{v}^{k}-\check{\bm{v}}^{k}\|_{\bm{G}}^{2},$ (S1.24)
$\|\bm{M}(\bm{v}^{k+1}-\check{\bm{v}}^{k+1})\|_{\bm{H}}^{2}\leq\|\bm{M}(\bm{v}^{k}-\check{\bm{v}}^{k})\|_{\bm{H}}^{2}.$ (S1.25)
The inequality in (S1.24) implies that the sequence $\{\bm{v}^{k}\}$ exhibits Fejér monotonicity. When $\bm{G}\succ 0$, this sequence converges to a point $\bm{v}^{*}\in\Omega^{*}$ in the $\bm{H}$-norm. This key inequality is fundamental for establishing the global convergence of the sequence. The proof of the global convergence can be found in Theorem 3.2 of [39]. So far, the proof of the first property of Theorem 3 has been completed.
Since $\bm{M}(\bm{v}^{k+1}-\check{\bm{v}}^{k+1})=\bm{v}^{k+1}-\bm{v}^{k+2}$ and $\bm{M}(\bm{v}^{k}-\check{\bm{v}}^{k})=\bm{v}^{k}-\bm{v}^{k+1}$, the inequality in (S1.25) implies that
$\|\bm{v}^{k+1}-\bm{v}^{k+2}\|_{\bm{H}}^{2}\leq\|\bm{v}^{k}-\bm{v}^{k+1}\|_{\bm{H}}^{2}.$ (S1.26)
From the inequality in (S1.24), it follows from $\bm{G}\succ 0$, there is a constant $c_{0}>0$ such that
$\|\bm{v}^{k+1}-\bm{v}^{*}\|_{\bm{H}}^{2}\leq\|\bm{v}^{k}-\bm{v}^{*}\|_{\bm{H}}^{2}-c_{0}\|\bm{M}(\bm{v}^{k}-\check{\bm{v}}^{k})\|_{\bm{H}}^{2}=\|\bm{v}^{k}-\bm{v}^{*}\|_{\bm{H}}^{2}-c_{0}\|\bm{v}^{k}-\bm{v}^{k+1}\|_{\bm{H}}^{2}.$ (S1.27)
By summing up inequality in (S1.27) over $k=0,1,\cdots,K$, we obtain
@@
Combining (S1.26) leads to the conclusion that
$c_{0}(K+1)\|\bm{v}^{K}-\bm{v}^{K+1}\|_{\bm{H}}^{2}\leq\|\bm{v}^{0}-\bm{v}^{*}\|_{\bm{H}}^{2}.$ (S1.28)
@@ Appendix D Proof of Throrem 4
$\left\{\hat{\bm{\beta}}^{k},\hat{\bm{z}}^{k},\hat{\bm{u}}^{k}\right\}=\left\{\bar{\bm{\beta}}^{k},\bar{\bm{z}}^{k},\bar{\bm{u}}^{t}\right\}.$ (S1.29)
@@
$\hat{\bm{\beta}}^{k+1}\leftarrow\arg\min_{\bm{\beta}}\left\{\lambda\|\bm{w}\odot\bm{\beta}\|_{1}+\frac{\eta}{2}\|\bm{\beta}-\hat{\bm{\beta}}^{k}+\frac{\mu\bm{X}^{\top}(\bm{X}\hat{\bm{\beta}}^{k}-\hat{\bm{r}}^{k}-\hat{\bm{u}}^{k}/\mu)}{\eta}\|_{2}^{2}\right\}.$ (S1.30)
@@
$\bar{\bm{\beta}}^{k+1}\leftarrow\underset{\bm{\beta}}{\mathrm{argmin}}\left\{\lambda\left\lVert\bm{w}\odot\bm{\beta}\right\rVert_{1}+\frac{\eta}{2}\left\lVert\bm{\beta}-\bar{\bm{\beta}}^{k}+\sum_{g=1}^{G}\bm{\xi}_{g}^{k}\right\rVert_{2}^{2}\right\},$ (S1.31)
@@
$\sum_{g=1}^{G}\bm{\xi}_{g}^{k}=\frac{\mu\bm{X}^{\top}(\bm{X}\bar{\bm{\beta}}^{k}-\bar{\bm{r}}^{k}-\bar{\bm{u}}^{k}/\mu)}{\eta}.$ (S1.32)
According to the assumption in (S1.29), there is $\hat{\bm{\beta}}^{k}=\bar{\bm{\beta}}^{k}$, $\hat{\bm{r}}^{k}=\bar{\bm{r}}^{k}$ and $\hat{\bm{u}}^{k}=\bar{\bm{u}}^{k}$. It follows from (S1.30) and (S1.31) that
$\hat{\bm{\beta}}^{k+1}=\bar{\bm{\beta}}^{k+1}.$ (S1.33)
@@
$\hat{\bm{r}}^{k+1}\leftarrow\left[\Phi(\hat{\bm{r}}^{k})+\mu\bm{I}_{p}\right]^{-1}\left[\mu(\bm{X}\hat{\bm{\beta}}^{k+1}-\hat{\bm{u}}^{k}/\mu)+\Phi(\hat{\bm{r}}^{k})\hat{\bm{r}}^{k}-\phi^{{}^{\prime}}(\hat{\bm{r}}^{k})+\bm{S}\right].$ (S1.34)
$\hat{\bm{r}}^{k+1}$ can be expressed as $(\hat{\bm{r}}_{1}^{k+1})^{\top},(\hat{\bm{r}}_{2}^{k+1})^{\top},\ldots,(\hat{\bm{r}}_{G}^{k+1})^{\top}]$, and $\bm{S}$, $\hat{\bm{u}}^{k}$ and $\hat{\bm{r}}^{k}$ are divided in a similar manner. Drawing on the definition of $\Phi(\hat{\bm{r}}^{k})$, it becomes clear that the diagonal elements of its block-diagonal matrix are $\Phi(\hat{\bm{r}}_{g}^{k})$, with $g$ ranging from $1$ to $G$. Then, we have
$\hat{\bm{r}}_{g}^{k+1}\leftarrow\left[\Phi(\hat{\bm{r}}_{g}^{k})+\mu\bm{I}_{n_{g}}\right]^{-1}\left[\mu(\bm{X}\hat{\bm{\beta}}_{g}^{k+1}-\hat{\bm{u}}_{g}^{k}/\mu)+\Phi(\hat{\bm{r}}_{g}^{k})\hat{\bm{r}}_{g}^{k}-\phi^{{}^{\prime}}(\hat{\bm{r}}_{g}^{k})+\bm{S}_{g}\right],g=1,\dots,G.$ (S1.35)
@@
$\bar{\bm{r}}_{g}^{k+1}\leftarrow\left[\Phi(\bar{\bm{r}}_{g}^{k})+\mu\bm{I}_{n_{g}}\right]^{-1}\left[\mu(\bm{X}_{g}\bar{\bm{\beta}}^{k+1}-\bar{\bm{u}}_{g}^{k}/\mu)+\Phi(\bar{\bm{r}}_{g}^{k})\bar{\bm{r}}_{g}^{k}-\phi^{{}^{\prime}}(\bar{\bm{r}}_{g}^{k})+\bm{S}_{g}\right],g=1,\dots,G.$ (S1.36)
Again, according to the assumption in (S1.29) and (S1.33), there is $\hat{\bm{r}}_{g}^{k}=\bar{\bm{r}}_{g}^{k}$, $\hat{\bm{u}}_{g}^{k}=\bar{\bm{u}}_{g}^{k}$ and $\hat{\bm{\beta}}^{k+1}=\bar{\bm{\beta}}^{k+1}$. It follows from (S1.35) and (S1.36) that
$\hat{\bm{r}}_{g}^{k+1}=\bar{\bm{r}}_{g}^{k+1},g=1,\dots,G.$ (S1.37)
@@
It follows from (S1.29), (S1.33) and (S1.37) that
$\hat{\bm{u}}_{g}^{t+1}=\bar{\bm{u}}_{g}^{t+1},g=1,\dots,G.$ (S1.40)
@@
$\left\{\hat{\bm{\beta}}^{k},\hat{\bm{r}}^{k},\hat{\bm{u}}^{k}\right\}=\left\{\bar{\bm{\beta}}^{k},\bar{\bm{z}}^{k},\bar{\bm{u}}^{k}\right\},\quad\text{for all }k.$ (S1.42)